Equivalent Units of Production and Related Costs

The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process-Assembly Department
Bal., 8,000 units, 60% completed 23,360 To Finished Goods, 184,000 units ?
Direct materials, 188,000 units @ $1.90 357,200
Direct labor 236,100
Factory overhead 91,860
Bal., ? units, 25% completed ?

Determine the following:

a. The number of units in work in process inventory at the end of the period.
units

b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "0".

Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs
Whole
Units
Equivalent Units
Direct Materials
Equivalent Units
Conversion
Inventory in process, beginning
Started and completed
Transferred to finished goods
Inventory in process, ending
Total units

c. Costs per equivalent unit for direct materials and conversion. If required, round your answers to the nearest cent.

Costs Per Equivalent Unit
Direct Materials $
Conversion $

d. Cost of the units started and completed during the period.
$
Answer #1
a
Number of units in work in process inventory 12000 =8000+188000-184000
b
Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs
Whole Equivalent Units Equivalent Units
Units Direct Materials Conversion
Inventory in process, beginning 8000 0 3200 =8000*(1-60%)
Started and completed 176000 176000 176000 =184000-8000
Transferred to finished goods 184000 176000 179200
Inventory in process, ending 12000 12000 3000 =12000*25%
Total units 196000 188000 182200
c
Costs Per Equivalent Unit
Direct Materials 1.90 =357200/188000
Conversion 1.80 =(236100+91860)/182200
d
Units started and completed 176000
X Costs per equivalent unit 3.70 =1.9+1.8
Cost of the units started and completed 651200